A researcher reported the results of a telephone poll of 1000 adult Americans. The question posed of those who were surveyed was: "Should the federal tax on cigarettes be raised to pay for health care reform?" Of 600 non-smokers, 362 said yes. Of 400 smokers, 80 said yes. What are the results at alpha = .05, if you want to determine that the proportion of non-smokers who said yes is greater than the proportion of smokers who said yes?
1
Expert's answer
2020-04-06T15:55:10-0400

"p_1=\frac{362}{600}=0.6033,\;\;p_2=\frac{80}{400}=0.2,\;\; p_0=\frac{362+80}{600+400}=0.442."

Null hypothesis "H_0: p_1=p_2."

Alternative hypothesis "H_a:p_1>p_2 ."

Test statistic: "z=\frac{p_1-p_2}{\sqrt{p_0(1-p_0)(\frac{1}{n_1}+\frac{1}{1/n_2}})}=
\frac{0.6033-0.2}{\sqrt{0442(1-0.442)(\frac{1}{600}+\frac{1}{400}})}=12.58."

P-value: "p<0.0001."

Since the P-value is less than 0.05, reject the null hypothesis.

There is a sufficient evidence that the proportion of non-smokers who said yes is greater than the proportion of smokers who said yes.